How much should the Earth be compressed without losing any material in order for it to become a blackhole. (note : The mass of Earth is 5.9 x 1024 kg and the gravitational constant is 6.67 x 10-11 N m2 kg-2 (Hint: Use Schwarzschild radius)
1
Expert's answer
2020-08-31T12:47:10-0400

Explanations & Calculations


  • Any object to attain the blackhole status is defined by a concept of Schwarzschild radius.
  • Any object reduced beyond volume of a sphere of a Schwarzschild radius, is expected to behave as a blackhole.
  • Schwarzschild radius is given by "r_s = \frac{2GM}{c^2}" where M & c are the mass of the object and the speed of light respectively.
  • Therefore,

"\qquad\qquad
\begin{aligned}
\small r_s &= \small \frac{2( 6.67\times10^{-11}Nm^2kg^{-2})
( 5.9\times 10^{24}kg)}{(3\times 10^8ms^{-1})^2}\\
\small &= \small 8.745\times10^{-3}m=8.745 mm
\end{aligned}"

  • Therefore, the Earth should be compressed down to a sphere of radius 8.745 mm.